consider a particle initially moving with 5 m s-1 starts decelerating at a constant rate of 2 m s -2. find the distance travelled in the third second.

Grade:11

1 Answers

Khimraj
5 years ago
velocity after 2 sec = 5 – 2*2 = 1m/s
so distance covered in next 1/2 sec = 1*1 – ½*2*(1/2)2 = ¾ m
now its velocity become zero and reverse the direction
then distance in next ½ sec = ½*2*(1/2)2 = ¼ M
SO distance in 3rd sec = ¾ + ¼ = 1m.